A dice is thrown two times. What is the probability that the sum Is neither 8 nor 9?


1
Expert's answer
2020-09-21T14:48:40-0400

If a dice is thrown 2 times then the total number of outcomes are as under:-


6C1*6C1 = 36


which can be represented as under-




The number of outcomes where the sum of the number on the dice is neither 8 nor 9 is marked in yellow in the above table = 27


Hence the probability P of getting an outcome from any of the yellow marked numbers


P = 27/36 = 0.75